Equipment and Calibration History
Page: [1] 2 3 4 5 6 7 8 9 10 11 12 13 14 15 16 17 18 19 20 21 22 23 24 25 26 27 28 29 30 31 32 33 34 35 36 37 38 39 40 41 42 43 44 45 46 47 48 49 50 51 52 53 54 55 56 57 58 59 60 61 62 63 64 65 66 67 68 69 70Ser No. | Date | Loc From | Title | Loc To | Notes |
---|---|---|---|---|---|
LI-7500A 75H-2514 | 2025-05-09 | Lab | Calibration | Lab | Last calibration: 2025-01-28 AGC=99.6 SDM=7 DAC1=- DAC2=- AIU sn AIU-2652, Presure A0=59.516, A1=15.503 Ambient P=100.8kPa, Tair=16.9C, Tdew=8.9C Zero Air at 100psi, AMP CA03923 488.75 at 1200 Tdewspan=13.0 CO2 zero 0.9083 -> 0.9091 Offset: 3.4ppm CO2 span 0.9983 -> 1.000 Offset: -0.95ppm H2O zero 0.8498 -> 0.8479 Offset: -12.7ppm H2O span 0.97 -> 0.9976 Offset: -0.57ppm Notes: The CO2 span was finicky. Tried a couple of small offsets from span gas value but there is anecdotal evidence that coefficient near 1.000 are difficult. Overall okay. |
LI-7500A 75H-2667 | 2025-05-08 | Lab | Calibration | Lab | Last calibration: 2025-01-07 AGC=99.5 SDM=7 DAC1=- DAC2=- AIU sn AIU-2652, Presure A0=59.516, A1=15.503 Ambient P=101.0kPa, Tair=19.1C, Tdew=9.8C Zero Air at 100psi, AMP CA03923 488.75 at 1200 Tdewspan=13.0 CO2 zero 0.941 -> 0.9418 Offset: ppm CO2 span 1.0165 -> 1.0218 Offset: ppm H2O zero 0.8454 -> 0.8421 Offset: ppm H2O span 1.0393 -> 1.0332 Offset: ppm Notes: I started to calibrate this head yesterday but the H2O zero offset was large ~-30 and opposite for the third time. I decided to swap the internal chemicals - they were hard and needed a screw driver to pry out of the bottles. I tried to flush the body with zero air but the zeros were still high ~18ppm CO2. I let it run overnight. The chopper is a little noisy and I had to use 0.1ppm less to get the CO2 span to read right. |
LI-7550 AIU-2023 | 2025-05-07 | Lab none | Loaned | Loaned none | Loaned to Delfina Grinspan for use this summer in vineyards in Sonoma and Napa. |
LI-7500RS 75H-1653 | 2025-05-07 | Lab none | Loaned | Loaned none | Loaned to Delfina Grinspan for use this summer in vineyards in Sonoma and Napa. |
HFP01 003072 | 2025-05-07 | Lab none | Loaned | Loaned none | Loaned to Delfina Grinspan for use this summer in vineyards in Sonoma and Napa. |
NR-LITE 041556 | 2025-05-07 | Lab none | Loaned | Loaned none | Loaned to Delfina Grinspan for use this summer in vineyards in Sonoma and Napa. |
WM 1590 153804 | 2025-05-07 | Lab none | Loaned | Loaned none | Loaned to Delfina Grinspan for use this summer in vineyards in Sonoma and Napa. |
LI-7500A 0035 | 2025-05-02 | Lab none | Loaned | Loaned none | Loaned to Patty to swap out the head on Eden Landing for testing and calibration. |
LI-7500A 0042 | 2025-05-07 | Lab | Calibration | Lab | Last calibration: 2025-01-07 AGC=102.5 SDM=7 DAC1=- DAC2=- AIU sn AIU-2652, Presure A0=59.516, A1=15.503 Ambient P=100.9kPa, Tair=20.1C, Tdew=8.7C Zero Air at 200psi, AMP CA03923 488.75 at 1200 Tdewspan=13.0 CO2 zero 0.9479 -> 0.9484 Offset: 2.1ppm CO2 span 1.0101 -> 1.0196 Offset: -5.95ppm H2O zero 0.8463 -> 0.8434 Offset: -17.7ppm H2O span 1.0236 -> 1.0342 Offset: -0.21ppm Notes: Some larger than usual offsets in co2 span and h2o zero but probably okay. |
LI-7500 0976 | 2025-05-02 | Lab none | Upgrade | Lab none | I upgraded this unit from 7500 to 7500A by replacing the bottom interface plate. Following the instructions I also adjusted the AGC pot and changed the signal strength coefficients. I also swapped the internal chemical - they were pretty hard and needs prying with a screw driver to get them out of the bottles. I'll let the head run over the weekend before I calibrate it. |
LI-7500 0976 | 2025-05-07 | Lab | Calibration | Lab | Last calibration: 2023-07-21 AGC=100.8 SDM=7 DAC1=- DAC2=- AIU sn AIU-2652, Presure A0=59.516, A1=15.503 Ambient P=100.8kPa, Tair=19.4C, Tdew=9.77C Zero Air at 200psi, AMP FB04304 401.79 at 300 Tdewspan=13.0 CO2 zero 0.8604 -> 0.8566 Offset: -5.3ppm CO2 span 1.0091 -> 1.0167 Offset: -8.59ppm H2O zero 0.6284 -> 0.6226 Offset: -3.5ppm H2O span 1.0207 -> 1.0272 Offset: 12.32ppm Notes: First cal after upgrade to 7500A and chemical swap. I reset signal strength from 76.1 to 100.0. H2O span was less stable than I would like. The original config file downloaded from Licor was use with co2 zero = 0.8579, co2 span = 1.000, h2o zero = 0.6231, h2o span = 0.9943. SD coefs set to zero. |
LI-7500A 0035 | 2025-05-02 | Lab | Calibration | Lab | Last calibration: 2025-01-07 AGC=100.1 SDM=7 DAC1=- DAC2=- AIU sn AIU-2652, Presure A0=59.516, A1=15.503 Ambient P=100.6kPa, Tair=16.4C, Tdew=9.94C Zero Air at 200psi, AMP FB04304 401.79 at 300 Tdewspan=13.0 CO2 zero 0.9373 -> 0.9374 Offset: 0.38ppm CO2 span 1.017 -> 1.0298 Offset: -6.39ppm H2O zero 0.8979 -> 0.9001 Offset: 13.3ppm H2O span 1.0284 -> 1.0280 Offset: 0.01ppm Notes: Used 401.69 to make span match. kind of large offset in CO2 span and H2O zero. |
PAR-LITE 030382 | 2025-04-29 | Lab none | Installed | Vaira PAR Sensor Reflected Cal | Installed near the existing reflected PAR sensor fro cross check. |
LI-7500A 0073 | 2025-04-29 | Lab none | Installed | Vaira Gas Analyzer CO2/H2O | Installed at Vaira to replace sn 0035 for calibration. |
LI-7500A 0035 | 2025-04-29 | Vaira Gas Analyzer CO2/H2O | Removed | Lab none | Removed from Vaira for calibration. Replaced by sn 0073. |
LI-7500A 0418 | 2025-04-29 | Lab none | Installed | Tonzi Gas Analyzer CO2/H2O Floor | Installed at Tonzi Floor to replace sn 0042 for calibration. |
LI-7500A 0042 | 2025-04-29 | Tonzi Gas Analyzer CO2/H2O Floor | Removed | Lab none | Removed from Tonzi Floor for calibration. Replaced by sn 0418. |
LI-7500A 0065 | 2025-04-29 | Lab none | Installed | Tonzi-Tower Gas Analyzer CO2/H2O | Installed at Tonzi Tower to replace sn 75H-2667 for calibration. |
LI-7500A 75H-2667 | 2025-04-29 | Tonzi-Tower Gas Analyzer CO2/H2O | Removed | Lab none | Removed from Tonzi-Tower for calibration. Replaced by sn 0065. |
LI-7700 TG1-0216 | 2025-04-17 | Lab none | Installed | Tonzi Gas Analyzer CH4 Floor | Installed at Tonzi Floor to replace sn TG1-0514 for calibration. |
LI-7700 TG1-0514 | 2025-04-17 | Tonzi Gas Analyzer CH4 Floor | Removed | Lab none | Removed from Tonzi Floor for calibration. Replaced by sn TG1-0216. |
LI-7500A 0065 | 2025-04-25 | Lab | Calibration | Lab | Last calibration: 2025-01-13 AGC=98.1 SDM=7 DAC1=- DAC2=- AIU sn AIU-2652, Presure A0=59.516, A1=15.503 Ambient P=100.3kPa, Tair=15.1C, Tdew=6.9C Zero Air at 200psi, AMP FB04304 401.79 at 750 Tdewspan=13.0 CO2 zero 0.8914 -> 0.8917 Offset: 1.75ppm CO2 span 1.0226 -> 1.0266 Offset: -1.99ppm H2O zero 0.7334 -> 0.7335 Offset: 0.5ppm H2O span 1.0237 -> 1.0285 Offset: -0.09ppm Notes: Had a SS of 110 in the field - not sure why much less in the lab |
LI-7700 TG1-0514 | 2025-04-25 | Lab | Calibration | Lab | Calibration Location: Lab Last calibration: 2024-10-07 RSSI=84.6%, Optical RH=10% Ambient P=100.37kPa, Tair=15.7C Zero Air at 200psi, AMP FB04304 1.9024 at 700 CH4 Zero Offset: -0.026ppm CH4 Span Offset: 0.022ppm Notes: Pressure took a long time to stabilize but eventually looked good matching weather apps and 7500. |
TEROS21 G20007096 | 2025-04-24 | Lab none | Tested | Lab none | Tested on a CR1000 in the lab - seems to be okay. Terminal SDI-12 interface worked okay and the sensor responded to water. Try reinstalling it in the field. |
TEROS21 G20007096 | 2024-12-03 | Tonzi Soil Water Potential C | Removed | Lab none | Removed from near OSU SP1 because of bad readings. No obvious damage. |
MiniCH4 38-576-25 | 2025-04-24 | Lab none | Tested | Lab none | I tested this sensor in the lab. The serial interface and temperature seem fine but the CH4 is reading -10.00. I downloaded all the data and erased the internal data file. I will try to get advice from Pro-Oceanus. |
MiniCH4 38-573-25 | 2025-04-24 | Lab none | Tested | Lab none | I tested this sensor in the lab. The serial interface and temperature seem fine but the CH4 is reading -10.00. I downloaded all the data and erased the internal data file. I will try to get advice from Pro-Oceanus. |
MiniCH4 38-573-25 | 2025-04-17 | Gilbert Tract CH4 Probe Tower | Removed | Lab none | Removed from float at GT boardwalk after animals chewed its cable, it also started to read -10.00 % CH4. |
MiniCH4 38-576-25 | 2025-04-16 | Mayberry CH4 Probe | Removed | Lab none | Removed from Mayberry due to reading -10.00 again. No problems with the mounting this time. |
HYDROS21 H21G00003238 | 2025-04-24 | Lab none | Tested | Lab none | Tested on a CR1000 in the lab. Seems like it is working fine. I think the failure in the field was do to the extension wires getting chewed or other external wiring issue. The sensor seems fine for reinstall. |
LI-7500A 75H-2176 | 2025-04-23 | Lab none | Installed | Mayberry Gas Analyzer CO2/H2O | Installed at Mayberry to replace sn 0065 for calibration. |
LI-7500A 0065 | 2025-04-23 | Mayberry Gas Analyzer CO2/H2O | Removed | Lab none | Removed from Mayberry for calibration. Replaced by sn 75H-2176. |
LI-7500RS 75H-1653 | 2025-04-21 | Lab | Calibration | Lab | Last calibration: 2023-09-08 AGC=101.3 SDM=1 DAC1=- DAC2=- AIU sn AIU-2023, Presure A0=59.172, A1=15.673 Ambient P=100.4kPa, Tair=19.7C, Tdew=11.0C Zero Air at 200psi, AMP FB04304 401.79 at 900 Tdewspan=16.0 CO2 zero 1.1420 -> 1.403 Offset: -3.7ppm CO2 span 0.9691 -> 0.9969 Offset: -14.39ppm H2O zero 1.0079 -> 1.0088 Offset: 4.1ppm H2O span 1.0189 -> 1.0211 Offset: -0.05ppm Notes: First cal after recieving from USGS with Rush Ranch stuff. Larger offsets not unexpected. We actually had this IRGA in 2011 from Franks as a loaner for Vaira. |
LI-7500A 75H-2176 | 2025-04-22 | Lab | Calibration | Lab | Last calibration: 2024-10-29 AGC=103.0 SDM=7 DAC1=- DAC2=- AIU sn AIU-2652, Presure A0=59.516, A1=15.503 Ambient P=100.3kPa, Tair=18.6C, Tdew=10.3C Zero Air at 200psi, AMP FB04304 401.79 at 800 Tdewspan=16.0 CO2 zero 0.8681 -> 0.8683 Offset: 0.90ppm CO2 span 1.0094 -> 1.0170 Offset: -3.79ppm H2O zero 0.8559 -> 0.8553 Offset: -4.0ppm H2O span 0.925 -> 0.9352 Offset: -0.23ppm Notes: Done with new 7550 box. All calibrations in this batch seem a little low on CO2 span. |
LI-7500A 0073 | 2025-04-22 | Lab | Calibration | Lab | Last calibration: 2024-10-29 AGC=104.0 SDM=7 DAC1=- DAC2=- AIU sn AIU-2652, Presure A0=59.516, A1=15.503 Ambient P=100.3kPa, Tair=18.7C, Tdew=10.3C Zero Air at 200psi, AMP FB04304 401.79 at 750 Tdewspan=16.0 CO2 zero 0.9156 -> 0.9159 Offset: 1.0ppm CO2 span 1.006 -> 1.0118 Offset: -2.89ppm H2O zero 0.8332 -> 0.8348 Offset: 8.8ppm H2O span 1.0191 -> 1.0282 Offset: -0.18ppm Notes: Done with new 7550 box. All the calibrations in this round seemed a little low on co2 span. |
WL-400-003 1202000485 | 2025-05-04 | Lab none | Installed | Rush Ranch Water Level Annex | Installed at the new Annex tower location but not wired yet. The well is two concentric pipes with gravel and filter screen and the bottom and sand and scree between the pipes. It's three feet deep with a foot of gravel on the bottom. |
LI-7700 TG1-0216 | 2025-04-15 | Lab | Calibration | Lab | Calibration Location: Lab Last calibration: 2024-08-16 RSSI=78.1%, Optical RH=4% Ambient P=100.6kPa, Tair=19.1C Ultra High Purity N2 at 550psi, AMP FB04304 1.9024 at 1000 CH4 Zero Offset: -0.031ppm CH4 Span Offset: 0.0076ppm Notes: Calibrated after returned from use at Eden Landing. Touched up scrapes on spars with nail polish |
CS450 20010770 | 2025-04-15 | Lab none | Tested | Lab none | I tested this sensor in the lab. Its power draw was hanging up a CR1000 or maybe the Lithium power pack. I tested it on the lab power supply and it was pulling 1.4A when it should be more like 80uA. The cable is pretty beat up with lots of tape patches, cracking and one abraded spot with the internal conductors visible. Not sure if it is worth cutting off the worst parts of the cable and trying it again. |
CS450 20010770 | 2024-12-19 | Bouldin Corn Water Level | Removed | Lab none | Removed from BC because it was shorting out all the 12V power. |
HYDROS21 H21G00003238 | 2025-04-08 | Gilbert Tract Water Level Boardwalk | Removed | Lab none | Disconnected that the junction box and brought in. This sensor stop working a couple of weeks ago. |
SL-610-SS 1408 | 2025-03-26 | Lab none | Installed | Rush Ranch Pyrgeometer Annex Outgoing | Installed on a Campbell tripod and moved out into the marsh south west of the main tower. Exact location still TBD |
Quantum Q40957 | 2025-03-26 | Lab none | Installed | Rush Ranch PAR Sensor Annex Outgoing | Installed on a Campbell tripod and moved out into the marsh south west of the main tower. Exact location still TBD |
CM11 996985 | 2025-03-26 | Lab none | Installed | Rush Ranch Pyranometer Annex Outgoing | Installed on a Campbell tripod and moved out into the marsh south west of the main tower. Exact location still TBD |
CS616 ag-tdr01 | 2025-03-26 | Lab none | Installed | Rush Ranch Soil Moisture Annex C | Installed on a Campbell tripod and moved out into the marsh south west of the main tower. Exact location still TBD |
CS655 67372 | 2025-03-26 | Lab none | Installed | Rush Ranch Soil Moisture Annex B | Installed on a Campbell tripod and moved out into the marsh south west of the main tower. Exact location still TBD |
CS655 67373 | 2025-03-26 | Lab none | Installed | Rush Ranch Soil Moisture Annex | Installed on a Campbell tripod and moved out into the marsh south west of the main tower. Exact location still TBD |
HP-127-1.0-1.3-71P PC4 | 2025-03-26 | Lab none | Installed | Rush Ranch Heat Flux Plate Annex C | Installed on a Campbell tripod and moved out into the marsh south west of the main tower. Exact location still TBD |
HP-127-1.0-1.3-71P PC3 | 2025-03-26 | Lab none | Installed | Rush Ranch Heaf Flux Plate Annex B | Installed on a Campbell tripod and moved out into the marsh south west of the main tower. Exact location still TBD |
HFP01 003047 | 2025-03-26 | Lab none | Installed | Rush Ranch Heaf Flux Plate Annex A | Installed on a Campbell tripod and moved out into the marsh south west of the main tower. Exact location still TBD |
HMP45A F1830005 | 2025-03-26 | Lab none | Installed | Rush Ranch RH & Temp Annex | Installed on a Campbell tripod and moved out into the marsh south west of the main tower. Exact location still TBD |
RF400 21520 | 2025-03-26 | Lab none | Installed | Rush Ranch Radio Annex | Installed on a Campbell tripod and moved out into the marsh south west of the main tower. Exact location still TBD |
RF400 21519 | 2025-03-26 | Lab none | Installed | Rush Ranch Radio | Installed on a Campbell tripod and moved out into the marsh south west of the main tower. Exact location still TBD |
AM25T 6768 | 2025-03-26 | Lab none | Installed | Rush Ranch Multiplexer Annex | Installed on a Campbell tripod and moved out into the marsh south west of the main tower. Exact location still TBD |
CR1000 45312 | 2025-03-26 | Lab none | Installed | Rush Ranch Data Logger Annex | Installed on a Campbell tripod and moved out into the marsh south west of the main tower. Exact location still TBD |
MiniCH4 38-576-25 | 2025-03-26 | Lab none | Installed | Mayberry CH4 Probe | I reinstalled the miniCH4 on the northeast leg of the scaffolding. |
MB7389-100 68574-3 | 2025-03-26 | Lab none | Installed | Hill Slough Water Level | Installed at Hill slough in a slightly higher location off the boardwalk post. This replaces sn 76294-1 that got wet, corroded and failed. |
MB7389-100 76294-1 | 2025-03-26 | Hill Slough Water Level | Removed | Lab none | Removed from Hill Slough because it got wet and started to corrode - bad readings. |
MB7389-100 68574-2 | 2025-03-24 | Lab none | Tested | Disposed none | I tested this sensor in the lab. It is dead. It was powered with 12V instead of 5V in the field and died. Put in e-waste. |
MB7389-100 68574-2 | 2025-01-15 | Lab none | Installed | Mayberry Range Finder water | Installed to replace the existing sensor. Hopefully this one will have better quality data. We might have killed it with 12V instead of 5V. |
URM06 URM06-002 | 2025-01-15 | Mayberry Range Finder water | Removed | Lab none | Removed from the tower for a newer maybe better one. The water level range data from this sensor is a bit noisy but useful for gap filling. |
MB7389-100 68574-4 | 2025-01-29 | Lab none | Installed | Mayberry Range Finder water | Installed to replace the one killed with 12V instead of 5V last time. |
MB7389-100 68574-2 | 2025-01-29 | Mayberry Range Finder water | Removed | Lab none | Removed from the tower because it was killed with 12V instead of 5V last time. |
LI-720 CFS-00000 | 2025-03-20 | Lab none | Installed | Bouldin Alfalfa Flux Analyzer CO2 | Installed on the tower to replace the first 720. I think this is the first production run unit. |
LI-720 CW-CDR-064 | 2025-03-20 | Bouldin Alfalfa Flux Analyzer CO2 | Removed | Lab none | Removed from the tower at the request of Licor - maybe incompatibility of firmware and interface software |
MiniCH4 38-572-25 | 2025-03-13 | Lab none | Tested | Lab none | I ran this sensor in continuous mode for 10s of minutes and it never stopped reading -10 for CH4. I removed the filter and didn't seen any obvious signs in water getting in. Need to contact ProOcenanus. |
MiniCH4 38-576-25 | 2025-03-13 | Lab none | Tested | Lab none | Tested in the Lab. Seems okay. The CH4 was reading near zero 0.02% and I could make it climb by breathing into the filter. |
MiniCH4 38-574-25 | 2025-03-07 | Lab none | Installed | Gilbert Tract CH4 Probe Channel | Installed at the GT channel to replace sn 38-576-25 that has been reading -10% CH4 |
MiniCH4 38-576-25 | 2025-03-07 | Gilbert Tract CH4 Probe Channel | Removed | Lab none | This one started reading -10% CH4 for no apparent reason. It was removed for testing and replaced by sn 38-574-25 |
MiniCH4 38-572-25 | 2025-03-07 | Mayberry CH4 Probe | Removed | Lab none | This sensor had fallen out of the cable ties and its sensor may have been in the sediment. The CH4 readings steadily rose past 4% but it was dropping out before climbing. Removed for testing. |
eosFD FD20190027 | 2025-03-11 | Tonzi-SoilCO2 Forced Diffusion Tree 1 | Moved | Tonzi-SoilCO2 Forced Diffusion Open | Move back to the Open location after substituting for the broken FD in the tree1 location |
PRB-K Type #8 000000402165 | 2025-03-11 | Tonzi Soil Moisture Probe Under | Removed | Lab none | Removed from the Under location near the Floor tower due to smashed and broken pins |
eosFD FD20180026 | 2025-03-11 | Vaira Forced Diffusion West | Moved | Tonzi-SoilCO2 Forced Diffusion Tree 1 | Removed from the west location at Varia to replace the tree1 chamber at Tonzi that died |
RPI0W RPI0W-005 | 2025-03-12 | Lab none | Notes | Lab none | This pi is missing the clip for its camera cable. That happened a while ago. We've made it work so far by stuffing a small piece of thick paper (e.g. manila folder) to keep the camera cable secure in its port. |
MiniCH4 38-572-25 | 2025-01-15 | Lab none | Installed | Mayberry CH4 Probe | Installed on the northeast scaffolding foot at a fixed height. Interface box in the data logger enclosure. |
BB-Fab BB-GT | 2025-02-18 | Gilbert Tract none | Gilbert Tract none | Two AGM group 27 batteries were stolen out of the plastic battery box probably sometime on the evening of 02/18. System without power for about two weeks. | |
LI-7700 TG1-0216 | 2025-02-28 | Loaned none | Returned | Lab none | Returned to the lab from Patty. It came back with at least the teal ethernet cable. |
EXO fDOM 22C100780 | 2025-02-27 | Lab none | Calibration | Lab none | fDOM (QSU and RFU), Last cal: 2025-02-27 EXO sn 23H107001 EXO/Sensor Firmware: 1.0.86/3.0.0 ,Standard,PreCal,PostCal,RawValue,Temp Pt1,0.00 QSU,-0.09 QSU,0.01 QSU,0.00 QSU,23.418 °C Pt2,302.7 QSU,307.42 QSU,302.70 QSU,0.00 QSU,18.463 °C ,Standard,PreCal,PostCal,RawValue,Temp Pt1,0.00 RFU,0.02 RFU,0.00 RFU,0.00 RFU,23.198 °C Pt2,100.90 RFU,100.60 RFU,100.90 RFU,0.00 RFU,19.207 °C Note: Made fresh calibration solution yesterday using QSE solution from 2024-07-20 but then forgot to update temperature compensated standard value. Stored solution in fridge overnight and let warm up to room temp (3hrs) today before calibration |
EXO pH-ORP 24G102326 | 2025-02-27 | Lab none | Calibration | Lab none | ORP (mV), Last Cal: 2024-10-22 EXO sn 23H107001 EXO/Sensor Firmware 1.0.86/3.0.0 ,Standard,PreCal,PostCal,RawValue,Temp Pt1,227.0 mV,218.3 mV,227.0 mV,0.0 mV,22.018 °C Note: Old Zobell solution (mixed 2022-09-23) only about 0.4mV different from fresh Zobell solution mixed today and stored in small brown bottle. Old Zobell solution stored in its original large brown bottle. |
EXO pH-ORP 24G102326 | 2025-02-27 | Lab none | Calibration | Lab none | pH, Last cal: 2024-10-22 EXO sn 23H107001 EXO/Sensor Firmware: 1.0.86/3.0.0 ,Standard,PreCal,PostCal,RawValue,Temp Pt1,7.01 pH,7.11 pH,7.01 pH,-18.6 mV,20.021 °C Pt2,10.06 pH,10.18 pH,10.06 pH,-192.5 mV,19.981 °C Pt3,4.00 pH,4.09 pH,4.00 pH,153.2 mV,20.023 °C 4-7Slope,4-7mVPerDecade,7-10Slope,7-10mVPerDecade 171.5 mV,56.98 mV,174.1 mV,57.09 Note: Using standards from 2023-12-04 |
EXO Wiped Cond/Temp 24F100050 | 2025-02-27 | Lab none | Calibration | Lab none | Sp Cond (uS/cm),Last cal: 2024-10-22 EXO sn 23H107001 EXO/Sensor Firmware 1.0.86/3.0.5 CellConstant 0.47 ,Standard,PreCal,PostCal,RawValue,Temp Pt1,10000.0 µS/cm, 10035.5 µS/cm, 10000.0 µS/cm, 0.0 µS/cm,17.473 °C Note: Old 10,000 uS/cm solution from 2022 had some dirt in it and read about 200uS/cm higher than the fresh used for the calibration. We discarded the old solution. |
EXO Turbidity 24F102410 | 2025-02-27 | Lab none | Calibration | Lab none | Turbidity (FNU), Last cal: 2024-10-22 EXO sn 23H107001 EXO/Sensor Firmware: 1.0.86/3.0.5 ,Standard,PreCal,PostCal,RawValue,Temp Pt1,0.00 FNU,-0.04 FNU,0.00 FNU,0.00 FNU,23.311 °C Pt2,124.00 FNU,114.35 FNU,124.01 FNU,0.00 FNU,23.413 °C Pt3,1010.00 FNU, 931.74 FNU, 1009.99 FNU, 0.00 FNU, 23.782 °C Note: Much larger offset for 1010 FNU point that we have seen before, brand new cal solutions |
EXO Central Wiper 24F104086 | 2025-02-18 | Gilbert Tract | Removed | Lab none | Removed from the GT channel and brought back to the lab for routine calibration |
EXO fDOM 22C100780 | 2025-02-18 | Gilbert Tract fDOM | Removed | Lab none | Removed from the GT channel and brought back to the lab for routine calibration |
EXO pH-ORP 24G102326 | 2025-02-18 | Gilbert Tract pH ORP | Removed | Lab none | Removed from the GT channel and brought back to the lab for routine calibration |
EXO Turbidity 24F102410 | 2025-02-18 | Gilbert Tract Turbidity | Removed | Lab none | Removed from the GT channel and brought back to the lab for routine calibration |
EXO Wiped Cond/Temp 24F100050 | 2025-02-18 | Gilbert Tract Conductivity Sensor Channel | Removed | Lab none | Removed from the GT channel and brought back to the lab for routine calibration |
EXO3S 23H107001 | 2025-02-18 | Gilbert Tract EXO | Removed | Lab none | Removed from the GT channel and brought back to the lab for routine calibration |
eosFD FD20190027 | 2025-02-21 | Tonzi-SoilCO2 Forced Diffusion Open | Moved | Tonzi-SoilCO2 Forced Diffusion Tree 1 | Moved from the open location to the tree1 location to replace FD20200001 that had bad CO2 readings |
eosFD FD20200001 | 2025-02-21 | Tonzi-SoilCO2 Forced Diffusion Tree 1 | Removed | Lab none | Removed from the Tree1 location at Tonzi due to bad CO2 readings. |
LI-7700 TG1-0211 | 2025-02-19 | Mayberry | Calibration | Mayberry | Calibration Location: Mayberry Last calibration: 2024-08-07 RSSI=71%, Optical RH=12% Ambient P=102.0kPa, Tair=16.0C Zero Air at 450psi, AMP CA03923 3.3865 at 1500 CH4 Zero Offset: -0.042ppm CH4 Span Offset: -0.015ppm Notes: Slight breeze |
LI-7700 TG1-0223 | 2025-02-18 | Gilbert Tract | Calibration | Gilbert Tract | Calibration Location: Gilbert Tract Last calibration: 2024-08-07 RSSI=83%, Optical RH=12% Ambient P=101.5kPa, Tair=17.5C Zero Air at 500psi, AMP CA03923 3.3865 at 1400 CH4 Zero Offset: 0.014ppm CH4 Span Offset: 0.0095ppm Notes: Still day |
LI-7700 TG1-0357 | 2025-02-19 | Bouldin Corn | Calibration | Bouldin Corn | Calibration Location: Bouldin Corn Last calibration: 2024-08-07 RSSI=74%, Optical RH=23% Ambient P=102.1kPa, Tair=14.2C Zero Air at 350psi, AMP CA03923 3.3865 at 1350 CH4 Zero Offset: 0.024ppm CH4 Span Offset: 0.0085ppm Notes: After a rain shower, breeze starting to pick up |
LI-7700 TG1-0423 | 2025-02-19 | East End | Calibration | East End | Calibration Location: East End Last calibration: 2024-08-07 RSSI=76%, Optical RH=7% Ambient P=102.2kPa, Tair=14.5C Zero Air at 450psi, AMP CA03923 3.3865 at 1300 CH4 Zero Offset: 0.01ppm CH4 Span Offset: -0.0135ppm Notes: Starting to rain but hood kept mirrors clean during calibration |
LI-7700 TG1-0155 | 2025-02-18 | Hill Slough | Calibration | Hill Slough | Calibration Location: Hill Slough Last calibration: 2024-08-07 RSSI=85%, Optical RH=5% Ambient P=102.0kPa, Tair=15.0C Zero Air at 500psi, AMP CA03923 3.3865 at 1500 CH4 Zero Offset: 0.022ppm CH4 Span Offset: -0.0015ppm Notes: Only a slight breeze |
MiniCH4 38-576-25 | 2025-02-18 | Lab none | Installed | Gilbert Tract CH4 Probe Channel | Installed in the GT channel, co-located with the ADCP, EXO, and channel miniDOT. It has a cable running up the ADCP railing to the enclosure for power/comms. |
LI-7500A 0041 | 2025-02-18 | Lab none | Installed | Rush Ranch Gas Analyzer CO2/H2O | Installed at Rush Ranch to replace sn 0418 for calibration. |
LI-7500A 0418 | 2025-02-18 | Rush Ranch Gas Analyzer CO2/H2O | Removed | Lab none | Removed from Rush Ranch for calibration. Replaced by sn 0041. |
LI-720 CW-CDR-064 | 2025-01-30 | Lab none | Installed | Bouldin Alfalfa Flux Analyzer CO2 | Installed in the same mount as the LI-710 recording on a CR6 |
NR-LITE 990354 | 2025-02-14 | Loaned none | Received | Lab none | Received back from Patty Oikawa. She ended up fixing her chewed cables and didn't use this one. |
miniDOT-usb 180871 | 2025-02-13 | Lab none | Loaned | Loaned none | Loaned to Patty Oikawa for use at the Eden Landing lateral flux station. |